OpenStudy (anonymous):

What is the magnitude and direction of line GH if G(3, -2) and H(6, 4)?

OpenStudy (anonymous):

A. magnitude: 6.71 units; direction 63.43° B. magnitude: 6.71 units: direction: 26.57° C. magnitude: 3.61 units; direction: 26.57° D. magnitude: 3.61 units: direction: 63.43°

OpenStudy (anonymous):

please help D: I dont understand how to do this at all

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

For magnitude, use the distance formula

OpenStudy (anonymous):

that's it?

ganeshie8 (ganeshie8):

that's it for magnitude :)

ganeshie8 (ganeshie8):

cuz magnitude is just the distance between two points

OpenStudy (anonymous):

well I did the distance formula and I got 6.7082

ganeshie8 (ganeshie8):

good job !

OpenStudy (anonymous):

what about the direction?

ganeshie8 (ganeshie8):

for direction, u need to find the \(\text{slope}\) first

OpenStudy (anonymous):

do I use the slope formula?

ganeshie8 (ganeshie8):

Yes !

OpenStudy (anonymous):

when I used the slope formula I got 2? is that right or did I make a mistake?

ganeshie8 (ganeshie8):

thats the correct slope

ganeshie8 (ganeshie8):

direction = \(\tan^{-1}(2)\)

OpenStudy (anonymous):

I got 63.43! omg thank you so much you are life saver!!! so the answer is A, right?

ganeshie8 (ganeshie8):

thats correct ! np :)

OpenStudy (anonymous):

thanks for the medal too c:

ganeshie8 (ganeshie8):

:)

OpenStudy (arnavguddu):

try these formulas..... its will be useful Say from A(x1, y1) to B(x2, y2).... dx = x2 - x1 dy = y2 - y1 Distance = sqrt( dx^2 + dy^2 ) Direction or angle = tan inverse (dy/dx)

OpenStudy (anonymous):

thank you! but I already got my answer. I'll use your method for future references though c:

OpenStudy (arnavguddu):

ok.... grt...and no probs